# Optimal Control of Discrete-Time Nonlinear Systems

### v1

@@ Abstract
This paper focuses on optimal control problem for a class of discrete-time nonlinear systems. In practical applications, computation time is a crucial consideration in solving optimal control problem, especially when meeting the real-time requirements of control systems. To address this challenge, this study proposes a novel framework based on the optimal control method. Firstly, the original optimal control problem is transformed into an equivalent optimization problem, which is resolved using the Pontryagin’s maximum principle from the perspective of optimal control. Furthermore, explicit formulas for computing both the gradient and hessian matrix of the cost function are proposed. A numerical algorithm with rapid convergence and high efficiency is also provided.
@@ I Introduction
Optimal control problem (OCP) is fundamentally an optimization problem (OP) focused on determining the control inputs for a dynamical system to minimize or maximize a cost function over time. It can be argued that many natural processes are essentially driven by optimization, reflecting the universal principle of optimality [1]. In a world increasingly shaped by optimization-based control, the growing diversity of applications demonstrates its essential role and wide-ranging impact across various fields. For example, the optimization-based control strategies have enabled bipedal, quadrupedal, and humanoid robots to walk reliably in standard environments, leading to their first practical deployments, such as Boston Dynamics’ Spot robot [2]. These advancements, along with the state-of-the-art performance of optimal control in various applications, emphasize the increasing importance of theoretical research in this field.
@@
To overcome the limitations of the aforementioned methods, [11] introduced an approach that combines nonlinear optimization with the maximum principle to improve the efficiency of solving the OCP. This method primarily employs FBDEs for gradient computation, replacing the traditional gradient computation in optimization. In [12], the hybrid gradient-based method was applied to simultaneously optimize unmanned aerial vehicle trajectories and control signals, minimizing mission completion time and energy consumption while ensuring smooth and practical implementation. Similarly, the control gains of a controller based on a fully actuated system approach were optimized using the hybrid method to ensure high-precision trajectory tracking while satisfying angular and actuator constraints [13]. Besides, the approach has also been extended to address the OCP of switched nonlinear systems [14, 15]. It should be noted that both the selection of optimization algorithms and gradient computation are crucial for computational efficiency, with each playing an indispensable role. However, most existing studies primarily focus on gradient computation, placing relatively less emphasis on the design of optimization algorithms.
In this paper, the OCP for discrete-time nonlinear systems is considered. A novel solution framework is introduced to address this problem. Unlike previous research that relies on existing optimization algorithms [11, 12, 13, 14, 15], this work presents a novel optimization approach from the perspective of optimal control. While it extends the findings of [16] and [17], which redefine the optimization problem and propose a superlinear convergent iterative algorithm, this study specifically addresses dynamic problems, focusing on the challenges associated with effectively computing the gradient and hessian matrix. To tackle these challenges, a new approach based on the maximum principle is introduced for the first time, transforming the solution process into solving two sets of FBDEs. This provides a unified computational approach for developing higher-order optimal control algorithms. The contributions of this paper can be summarized as follows:
@@
    A unified and explicit formula based on the maximum principle for computing the gradient and hessian matrix in OCP is proposed in this work, which reduces the computational burden. Furthermore, this method can be generalized to other second-order optimization algorithms.
The rest of this paper is organized as follows. In section II, the discrete-time nonlinear OCP is given. In section III, the control design and the iterative algorithm’s implementation and details are presented. Section IV concludes the work.
@@ III Main Results
In this section, in order to solve the OP (4), the problem is reformulated as a new OCP from an optimal control perspective, and a rapidly convergent iteration algorithm is proposed. Subsequently, the explicit formulas for computing the gradient and hessian matrix of the cost function are proposed by solving the corresponding FBDEs.
@@ III-A Optimal Control Design
$\underset{\upsilon_{[0,M-1]}}{\text{min}}J_{M}=\sum_{i=0}^{M-1}[J_{N}(x_{0},z_{i})+\frac{1}{2}\upsilon_{i}^{T}R\upsilon_{i}]+J_{N}(x_{0},z_{M}),$ (5)
@@
where $z_{i}\in\mathbb{R}^{m(N+1)}$ and $\upsilon_{i}\in\mathbb{R}^{m(N+1)}$ are the state and control variables, respectively, $z_{0}$ is the initial state value, $R$ is the positive definite matrix, and the cost function (5) is based on that of OP (4).
@@
$\gamma_{i}$ $=\gamma_{i+1}+\nabla J_{N}(x_{0},z_{i}),$ (7)
@@
where (7) is the costate equation with the terminal condition $\gamma_{M}=\nabla J_{N}(x_{0},z_{M})$, and (8) is the equilibrium condition. It is known that any solution to the OCP must satisfy the FBDEs. Therefore, the optimal control and state of the OCP (5)-(6) can be obtained by solving the FBDEs (6)-(8). Although FBDEs (6)-(8) can be numerically solved using the existing MSA, its convergence rate is relatively slow. To improve computational efficiency, an iterative numerical algorithm has been developed by combining the Taylor expansion with mathematical induction in [16, 17], as outlined below.
@@
$g_{i}$ $=(R+\nabla^{2}J_{N}(x_{0},z_{i}))^{-1}(\nabla J_{N}(x_{0},z_{i})+Rg_{i-1}),$ (10)
$g_{0}$ $=(R+\nabla^{2}J_{N}(x_{0},z_{i}))^{-1}\nabla J_{N}(x_{0},z_{i}),$ (11)
@@
In the iterative algorithm (9)-(11), the update step $g_{i}$ requires inner loop iterations starting from $g_{0}$. It should be noted that the algorithm faces considerable challenges in handling dynamic problem, i.e., the OCP (1)-(2), due to the computational complexity involved in evaluating $\nabla J_{N}$ and $\nabla^{2}J_{N}$. The following subsections provide the explicit iterative solution algorithms. For details on the convergence analysis and the superlinear convergence rate of the algorithm (9)-(11), please refer to [16, 17].
@@ III-B Computation of Gradient
From the OCP (1)-(2), it can be seen that the cost function $J_{N}$ is a complex function on the control variables $u_{k}$ to be optimized. To enhance computational efficiency, the gradient formula for the cost function $J_{N}$ is provided in this subsection.
@@
For the OP (4), the gradient of $J_{N}$ with respect to $z$ is given by
$\nabla J_{N}=\left[\frac{\partial H_{N}(x_{0},u_{0},\lambda_{1})}{\partial u_{0}},...,\frac{\partial H_{N}(x_{N},u_{N},\lambda_{N+1})}{\partial u_{N}}\right]^{T},$ (13)
where $x_{k}$ and $H_{N}$ are given by (2) and (12), and $\lambda_{k}$ is given by
$\lambda_{k}=\frac{\partial H_{N}(x_{k},u_{k},\lambda_{k+1})}{\partial x_{k}},$ (14)
@@
Proof: The proof can be found in [11] and is omitted here.
@@
Step 3: Compute the gradient of $J_{N}$ according to (13).
@@ III-C Computation of Hessian Matrix
In this subsection, the second-order derivative of the cost function $J_{N}$ is computed based on Lemma 1, and the explicit formula for hessian matrix computation will be presented.
@@
$G(x_{k},u_{k},\lambda_{k+1})=\frac{\partial H_{N}(x_{k},u_{k},\lambda_{k+1})}{\partial x_{k}},$
$L(x_{k},u_{k},\lambda_{k+1})=\frac{\partial H_{N}(x_{k},u_{k},\lambda_{k+1})}{\partial u_{k}}.$
@@
For the OP (4), the elements of the $i$th row of the hessian matrix of the cost function $J_{N}$ with respect to $z$ are given by
$\nabla^{2}J_{N}(i)=\left[\frac{\partial H_{0}^{i}}{\partial u_{0}},\frac{\partial H_{1}^{i}}{\partial u_{1}},...,\frac{\partial H_{N}^{i}}{\partial u_{N}}\right],$ (22)
where $H^{i}_{k}=H_{i}(x_{k},u_{k},\lambda_{k+1},\alpha_{k+1},\beta_{k})$, and $x_{k},\lambda_{k+1},\alpha_{k+1}$ and $\beta_{k}$ are given by (16), (17), (20) and (21), respectively.
@@
where $x_{k}^{*},\lambda_{k+1}^{*},\alpha_{k+1}^{*}$ and $\beta_{k}^{*}$ satisfy (16), (17), (20) and (21), respectively. When the given $u_{k}$ does note meet the equilibrium condition (29), the left side of equation (29) represents the $k$th element of the gradient, i.e., the (22) represents the $i$th row of the hessian matrix of the cost function $J_{N}$. The proof is complete.
@@
Step 3: Compute the elements of the $i$th row of the hessian matrix of $J_{N}$ according to (22).
@@
5:         Calculate $\nabla J=\nabla J_{N}(x_{t_{0}},z_{i})$;
6:         Calculate $\nabla^{2}J=\nabla^{2}J_{N}(x_{t_{0}},z_{i})$;
@@ III-D Implementation
IV Conclusion

### v2

@@ Abstract
This paper focuses on optimal control problem for a class of discrete-time nonlinear systems. In practical applications, computation time is a crucial consideration when solving nonlinear optimal control problems, especially under real-time constraints. While linearization methods are computationally efficient, their inherent low accuracy can compromise control precision and overall performance. To address this challenge, this study proposes a novel approach based on the optimal control method. Firstly, the original optimal control problem is transformed into an equivalent optimization problem, which is resolved using the Pontryagin’s maximum principle, and a superlinear convergence algorithm is presented. Furthermore, to improve computation efficiency, explicit formulas for computing both the gradient and hessian matrix of the cost function are proposed. Finally, the effectiveness of the proposed algorithm is validated through simulations and experiments on a linear quadratic regulator problem and an automatic guided vehicle trajectory tracking problem, demonstrating its ability for real-time online precise control.
@@ I Introduction
Optimal control problem (OCP) is fundamentally an optimization problem (OP) focused on determining the control inputs for a dynamical system to minimize or maximize a cost function over time. It can be argued that many natural processes are essentially driven by optimization, reflecting the universal principle of optimality [1]. In a world increasingly shaped by optimization-based control, the growing diversity of applications demonstrates its essential role and wide-ranging impact across various fields. For example, the optimization-based control strategies have enabled bipedal, quadrupedal, and humanoid robots to walk reliably in standard environments, leading to their first practical deployments, such as Boston Dynamics’ Spot robot [2]. It is important to note that most real-world systems exhibit inherent nonlinearities. While linearization is commonly employed in control system design for simplification, it often results in the loss of important nonlinear characteristics [3, 4]. As the demand for enhanced performance increases, it becomes increasingly critical to account for the nonlinear nature of the system in the design and analysis process.
@@
To overcome the limitations of the aforementioned methods, [13] introduced an approach that combines nonlinear optimization with the maximum principle to improve the efficiency of solving the OCP. This method primarily employs FBDEs for gradient computation, replacing the traditional gradient computation in optimization. In [14], the hybrid gradient-based method was applied to simultaneously optimize unmanned aerial vehicle trajectories and control signals, minimizing mission completion time and energy consumption while ensuring smooth and practical implementation. Similarly, the control gains of a controller based on a fully actuated system approach were optimized using the hybrid method to ensure high-precision trajectory tracking while satisfying angular and actuator constraints [15]. Besides, the approach has also been extended to address the OCP of switched nonlinear systems [16, 17]. It should be noted that both the selection of optimization algorithms and gradient computation are crucial for computational efficiency, with each playing an indispensable role. However, the majority of existing studies devote relatively less attention to the design of optimization algorithms. Furthermore, the current methods for gradient calculation are often not readily applicable to higher-order optimization algorithms.
In this paper, the OCP for discrete-time nonlinear systems is considered. Unlike previous research that relies on existing optimization algorithms [13, 14, 15, 16, 17], this work presents a novel optimization approach from the perspective of optimal control, featuring a superlinear convergence property. Following the work of [18] and [19], this study specifically addresses dynamic problems, with a particular focus on the challenges associated with effectively computing the gradient and hessian matrix. To tackle these challenges, a new approach based on the maximum principle is introduced for the first time, transforming the solution process into solving two sets of FBDEs. This provides a unified computational approach for developing second-order optimal control algorithms. The contributions of this paper can be summarized as follows:
@@
    A unified and explicit formula based on the maximum principle for computing the gradient and hessian matrix in OCP is proposed in this work, which reduces the computational burden. Furthermore, this method can be generalized to other second-order algorithms.
  • 3)
    The proposed optimal control algorithm is validated numerically and experimentally on linear quadratic regulator (LQR) and AGV systems, with results demonstrating its capability to achieve real-time online precise control computation.
The rest of this paper is organized as follows. In section II, the discrete-time nonlinear OCP is given. In section III, the control design and the iterative algorithm’s implementation and details are presented. Section IV presents the simulation and experimental validation for both LQR and AGV systems. Section V concludes the work.
@@ III Main Results
In this section, in order to solve the OP (4), the problem is reformulated as a new OCP from an optimal control perspective, and a rapidly convergent iteration algorithm is presented. Subsequently, the explicit formulas for computing the gradient and hessian matrix of the cost function are proposed by solving the corresponding FBDEs.
@@ III-A Optimal Control Design
$\underset{\upsilon_{[0,M-1]}}{\text{min}}J_{M}=\sum_{i=0}^{M-1}[J(x_{0},z_{i})+\frac{1}{2}\upsilon_{i}^{T}R\upsilon_{i}]+J(x_{0},z_{M}),$ (5)
@@
where $z_{i}\in\mathbb{R}^{m(N+1)}$ and $\upsilon_{i}\in\mathbb{R}^{m(N+1)}$ are the state and control variables, respectively, $z_{0}$ is the initial state value, $R$ is the positive definite matrix. It should be noted that the design of the cost function (5) is based on that of OP (4). At each time step, the controller $v_{i}$ minimizes the sum of costs over future steps, i.e., $\sum_{j=i+1}^{M}J(x_{0},z_{j})$, thereby accelerating the convergence of the optimization algorithm. Furthermore, by integrating control energy minimization into the loss function (5), the algorithm can achieve enhanced stability.
@@
$\gamma_{i}$ $=\gamma_{i+1}+\nabla J(x_{0},z_{i}),$ (7)
@@
where (7) is the costate equation with the terminal condition $\gamma_{M}=\nabla J(x_{0},z_{M})$, and (8) is the equilibrium condition. It is known that any solution to the OCP must satisfy the FBDEs. Therefore, the optimal control and state of the OCP (5)-(6) can be obtained by solving the FBDEs (6)-(8). It is evident from the FBDEs (6)-(8) that when the state $z_{i}$ in the system (6) stabilizes, the terminal state $z_{M}$ represents the optimal solution to the OP (4).
Although FBDEs (6)-(8) can be numerically solved using the existing MSA, its convergence rate is relatively slow. To improve computational efficiency, an iterative numerical algorithm has been developed by combining the Taylor expansion with mathematical induction in [18, 19], as outlined below.
@@
$g_{i}$ $=(R+\nabla^{2}J(x_{0},z_{i}))^{-1}(\nabla J(x_{0},z_{i})+Rg_{i-1}),$ (10)
$g_{0}$ $=(R+\nabla^{2}J(x_{0},z_{i}))^{-1}\nabla J(x_{0},z_{i}),$ (11)
@@
In the iterative algorithm (9)-(11), the update step $g_{i}$ requires inner loop iterations starting from $g_{0}$. It should be noted that the algorithm faces considerable challenges in handling dynamic problem, i.e., the OCP (1)-(2), due to the computational complexity involved in evaluating $\nabla J$ and $\nabla^{2}J$. The following subsections provide the explicit iterative solution algorithms. For details on the convergence analysis and the superlinear convergence rate of the algorithm (9)-(11), please refer to [18, 19].
@@ III-B Computation of Gradient
From the OCP (1)-(2), it can be seen that the cost function $J$ is a complex function on the control variables $u_{k}$ to be optimized. To enhance computational efficiency, the gradient formula for the cost function $J$ is provided in this subsection.
@@
For the OP (4), the gradient of $J$ with respect to $z$ is given by
$\nabla J=\left[\frac{\partial H(x_{0},u_{0},\lambda_{1})}{\partial u_{0}},...,\frac{\partial H(x_{N},u_{N},\lambda_{N+1})}{\partial u_{N}}\right]^{T},$ (13)
where $x_{k}$ and $H$ are given by (2) and (12), and $\lambda_{k}$ is given by
$\lambda_{k}=\frac{\partial H(x_{k},u_{k},\lambda_{k+1})}{\partial x_{k}},$ (14)
@@
Proof: For the proof, please refer to [20], which is omitted here.
@@
Step 3: Compute the gradient of $J$ according to (13).
@@ III-C Computation of Hessian Matrix
In this subsection, the second derivative of the cost function $J$ is computed based on Lemma 1, and the explicit formula for hessian matrix computation will be presented.
@@
$G(x_{k},u_{k},\lambda_{k+1})=\frac{\partial H(x_{k},u_{k},\lambda_{k+1})}{\partial x_{k}},$
$L(x_{k},u_{k},\lambda_{k+1})=\frac{\partial H(x_{k},u_{k},\lambda_{k+1})}{\partial u_{k}}.$
@@
For the OP (4), the elements of the $i$th row of the hessian matrix of the cost function $J$ with respect to $z$ are given by
$\nabla^{2}J(i)=\left[\frac{\partial H_{0}^{i}}{\partial u_{0}},\frac{\partial H_{1}^{i}}{\partial u_{1}},...,\frac{\partial H_{N}^{i}}{\partial u_{N}}\right],$ (22)
where $H^{i}_{k}=H_{i}(x_{k},u_{k},\lambda_{k+1},\alpha_{k+1},\beta_{k})$, and $x_{k},\lambda_{k+1},\beta_{k}$ and $\alpha_{k+1}$ are given by (16), (17), (20) and (21), respectively.
@@
where $x_{k}^{*},\lambda_{k+1}^{*},\beta_{k}^{*}$ and $\alpha_{k+1}^{*}$ satisfy (16), (17), (20) and (21), respectively. When the given $u_{k}$ does note meet the equilibrium condition (29), the left side of equation (29) represents the $k$th element of the gradient, i.e., the (22) represents the $i$th row of the hessian matrix of the cost function $J$. The proof is complete.
@@
Step 3: Compute the elements of the $i$th row of the hessian matrix of $J$ according to (22).
@@
Remark 3
The maximum principle is applied three times throughout this work. In the first application, it solves OCP (5)-(6) to develop the iterative optimization algorithm for OP (4). In the second, it is used to derive the gradient of the cost function $J$ with respect to $z$ by solving OCP (1)-(2). The final application involves solving OCP (15)-(17) to provide the explicit formula for the second-order derivative of the cost function $J$ with respect to $z$.
@@
5:         Calculate $\nabla J=\nabla J(x_{t_{0}},z_{i})$;
6:         Calculate $\nabla^{2}J=\nabla^{2}J(x_{t_{0}},z_{i})$;
@@ III-D Implementation
IV Simulation and experimental results
This section presents the simulation and experimental results to demonstrate the effectiveness of the proposed optimal control algorithm. First, the simulation validation of the LQR example is provided, followed by the simulation and experimental results of AGV trajectory tracking control.
IV-A Analytical solution verification: LQR
Consider the following linear discrete-time system:
$x_{k+1}=ax_{k}+bu_{k},$
where $x_{k}\in\mathbb{R}$ is the state variable with the given initial value $x_{0}$, $u_{k}\in\mathbb{R}$ the control variable, and $a$ and $b$ are the given constants. The associated quadratic cost function is given by
$J=\sum_{k=0}^{N-1}[qx_{k}^{2}+ru_{k}^{2}]+px_{N}^{2},$
where $q\geq 0$, $r>0$, and $p\geq 0$. The optimal control problem outlined above can be analytically solved by solving the discrete Riccati equation to obtain the optimal controllers. This solution serves as a benchmark to assess the accuracy of Algorithm 1 in determining the controllers.
In this example, we choose the parameters $a=1.8$, $b=0.9$, $q=1.0$, $r=3.0$, $p=3.0$, $N=15$, and $\bar{R}=0.1$. Besides, three independent simulations are conducted under different initial state value $x_{0}$, with Algorithm 1 running a single iteration of the outer loop. Fig. 1 illustrates the convergence of optimal state and control obtained by Algorithm 1 and the analytical solution. It is evident that the results of Algorithm 1 nearly coincide with the analytical solution, as validated through three simulations.
Figure 1: Comparison of the optimal state and control obtained from Algorithm 1 and the analytical solution under three different initial conditions: $x_{0}=1.0,2.0,3.0$. (a) Optimal state. (b) Optimal control.
IV-B Application: Trajectory tracking control of AGV
IV-B1 Problem Formulation
Consider the following kinematic model of AGV
$\begin{cases}\dot{x}_{t}=v_{t}\cdot{\rm cos}\theta_{t},\\ \dot{y}_{t}=v_{t}\cdot{\rm sin}\theta_{t},\\ \dot{\theta}_{t}=\omega_{t},\end{cases}$ (30)
where $t\in[0,t_{f}]$. The AGV’s planar motion is characterized by three state variables: $x$, $y$, and $\theta$. Here, $x$ and $y$ denote the position in the plane, while $\theta$ represents the AGV’s orientation. The control inputs are $v$ and $\omega$, which correspond to the linear and angular speeds of the AGV, respectively.
Figure 2: Comparison of trajectory tracking results and relevant parameters among Algorithm 1, QN, and SQP. (a) The trajectory tracking results of Algorithm 1, LMPC and QN. (b) The computation time for running 10 independent simulations with each of the three algorithms. (c) The number of iterations required for optimization at each sampling time $k$ for the three algorithms.
The system described by the equation (30) is discretized using the forward Euler method, resulting in the following difference equations:
$\begin{cases}x_{k+1}=x_{k}+\Delta\cdot v_{k}{\rm cos}\theta_{k},\\ y_{k+1}=y_{k}+\Delta\cdot v_{k}{\rm sin}\theta_{k},\\ \theta_{k+1}=\theta_{k}+\Delta\cdot\omega_{k},\end{cases}$ (31)
where $k=0,1,...,N-1$ represents the sampling times, and $\Delta$ is the time step. For convenience, let the state be denoted by $X_{k}=[x_{k},y_{k},\theta_{k}]^{T}$ and the control input by $U_{k}=[v_{k},\omega_{k}]^{T}$. Additionally, denote the reference state and control input by $X_{r}=[x_{r},y_{r},\theta_{r}]^{T}$ and $U_{r}=[v_{r},\omega_{r}]^{T}$, respectively. The error between the actual and reference values is given by $X_{e}=X_{k}-X_{r}$ and $U_{e}=U_{k}-U_{r}$. The associated cost function is defined as
$J=\sum_{k=0}^{N-1}(X_{e}^{T}\mathcal{Q}X_{e}+U_{e}^{T}\mathcal{R}U_{e}),$ (32)
where $\mathcal{Q}$ and $\mathcal{R}$ are the weighting matrix. The objective is to design an optimal controller $U_{k}$ to enable the AGV to track the reference trajectory.
IV-B2 Simulation Results
In this example, set the terminal sampling time $N=410$, the time step $\Delta=0.05$, the prediction horizon $N_{p}=10$, the AGV’s initial pose $X_{0}=[1.0,0.5,1.0]$, and the weight matrix $\mathcal{Q}={\text{diag}}[150,150,3]$, $\mathcal{R}={\text{diag}}[0.5,0.5]$, and $R=0.1I_{2N_{p}\times 2N_{p}}$. The simulations are conducted on a computer equipped with a 13th Gen Intel(R) Core(TM) i7-13620H processor and 16.0 GB of RAM, running Matlab R2020b.
The OCP (31)-(32) can be solved using linearization MPC (LMPC) and existing optimization algorithms such as gradient descent, quasi-newton (QN), and sequential quadratic programming (SQP) [3, 4, 5]. Here, QN and SQP are selected as benchmarks for Algorithm 1. Although SQP is typically used for constrained optimization, it is also efficient for unconstrained problems. For implementation, QN and SQP algorithms are executed using Matlab’s built-in solvers fminunc and fmincon, respectively. It should be noted that both QN and SQP can be implemented with or without gradient information. Additionally, for each optimization, the initial values of the parameters are set to a zero vector, and the stopping criterion is defined as the infinity norm of $\nabla J$ being less than $10^{-6}$.
Fig. 2 presents the tracking results for the circular trajectory along with a comparison of relevant metrics. In Fig. 2(a), QN algorithm is used to validate the optimal controller obtained by Algorithm 1. The results demonstrate that the optimal state trajectory produced by Algorithm 1 aligns perfectly with the QN results, confirming the success of Algorithm 1 in solving the trajectory tracking problem. Besides, compared to LMPC, Algorithm 1 achieves faster response and higher precision control. Fig. 2(b) compares the computation time required by the three algorithms during the tracking process. When using QN and SQP algorithms, providing gradient information significantly reduces computation time, with both algorithms showing similar performance under these conditions. In contrast, Algorithm 1 achieves nearly a 10-fold reduction in computation time compared to these two algorithms. Fig. 2(c) illustrates the number of iterations required to compute the optimal controller at each sampling time. It is clear that Algorithm 1 requires far fewer iterations than QN and SQP to reach the same level of accuracy.
Figure 3: Experimental platform.
IV-B3 Experimental Results
In the experiment, a closed indoor environment is selected. The AGV, as illustrated in Fig. 3, is primarily equipped with: 1) a 2D LiDAR for localization; 2) a Jetson Nano B01 mini PC to run Algorithm 1; and 3) a dual DC motor drive module AT8236 to control the motors.
Fig. 4(a) illustrates an indoor environmental map constructed using simultaneous localization and mapping (SLAM) technology, where the gray area denotes the accessible region for the robot. The blue line represents the tracking trajectory, with the starting point located at the top-left corner of the map and the endpoint at the bottom-left corner. The validation process of the algorithm is carried out on the robot operating system (ROS) platform, with all the code written in C++. During the tracking procedure, we set up remote communication between a virtual machine running Ubuntu 20.04 and the vehicle to manage the AGV’s starting, stopping, and data collection processes.
Fig. 4(b) and 4(c) illustrate the positional changes and error characteristics during the tracking process. The initial position of the vehicle is approximately 0.08 meters away from the starting point of the trajectory. At time points 12 seconds and 22 seconds, the vehicle exhibits slight fluctuations while turning. Apart from these three instances, the tracking error at other positions remains below 0.02 meters. Fig. 4(d) and 4(e) depict the variations in controller errors during the tracking process. Initially, the vehicle undergoes an acceleration phase to realign with the desired tracking trajectory, resulting in relatively large errors in linear and angular velocities during this phase. In the following stages, the vehicle keeps its speed close to the target values. There are slight increases in angular velocities during turns, while the variations at other positions remain relatively stable.
Figure 4: The trajectory tracking results of AGV in indoor environment. (a) Map of the environment and the desired tracking trajectory. (b) AGV’s tracking performance. (c) Trajectory tracking error. (d) Linear velocity $u$ error. (e) Angular velocity $\omega$ error.
V Conclusion